# Distributed quantum error correction based on hyperbolic Floquet codes

### v1

@@ V Conclusion
We have shown that hyperbolic Floquet codes are a viable option for running quantum error-correcting codes over a network of many small quantum processors with near-term qubit numbers and error rates. These codes have efficient encoding rates and low weight stabilisers, making them tolerant to higher error rates between devices. Results from our numerical simulations show that distributed hyperbolic codes can tolerate non-local noise of up to 1.3% without entanglement distillation. This shows how the codes can tolerate non-local noise up to an magnitude higher than the threshold for local noise. For a phenomenological noise model with feasible local ($10^{-4}$) and non-local ($5\times 10^{-3}$) error rates, we find that hyperbolic Floquet codes achieve good error suppression and feasible resource requirements for running a million quantum operations. From this we conclude that distributed error correction offers a viable, practically feasible, and efficient path for scaling quantum computation. This enables a complementary path to scale alongside efforts to improve the number of physical qubits or qubit fidelities per device.
One area of further work is understanding how to perform logical gates on hyperbolic Floquet codes. For useful fault-tolerant quantum computation, the logical gates should have a low execution complexity. It is known that logical operations can be implemented on planar Floquet codes through lattice surgery [58]. However, other techniques with lower overhead may exist, motivated by the availability of long-range operations in a distributed architecture involving photonic interconnects. Long-range operations might support approaches to universality that avoid the $O(d^{3})$ space-time volume scaling seen for universal gate sets in strictly local settings. In principle, universal quantum gates are possible through braiding Anyon excitations, although this is only discussed with respect to Fibonacci codes[59] or with 3-D extensions[60]. It is not yet known if these operations can extend to high rate codes, or how errors might propagate between logical gates. We also hope that an error-correction-aware graph partitioning process may improve on the partitioning scheme used in this work, possibly leading to improved error thresholds.

### v2

@@ V Conclusion
We have shown that hyperbolic Floquet codes are a viable option for running quantum error-correcting codes over a network of many small quantum processors with near-term qubit numbers and error rates. These codes have efficient encoding rates and low weight stabilisers, making them tolerant to higher error rates between devices. Results from our numerical simulations show that the distributed hyperbolic code H144-f3 can tolerate non-local noise of up to 3% without entanglement distillation. This shows how the codes can tolerate non-local noise up to an order of magnitude higher than the threshold for local noise. For a circuit noise model with feasible local ($99.97\%$) and non-local ($99\%$) error rates, we find that hyperbolic Floquet codes achieve good error suppression and feasible resource requirements for running a million quantum operations. From this we conclude that distributed error correction offers a viable, practically feasible, and efficient path for scaling quantum computation. This enables a complementary path to quantum utility alongside efforts to improve the number of physical qubits or qubit fidelities per device.
One area of further work is understanding how to perform logical gates on hyperbolic Floquet codes. For useful fault-tolerant quantum computation, the logical gates should have low execution complexity. It is known that logical operations can be implemented on planar Floquet codes through lattice surgery [59]. However, other techniques with lower overhead may exist, motivated by the availability of long-range operations in a distributed architecture involving photonic interconnects. Long-range operations might support approaches to universality that avoid the $O(d^{3})$ space-time volume scaling seen for universal gate sets in strictly local settings. In principle, universal quantum gates are possible through braiding Anyon excitations, although this is only discussed with respect to Fibonacci codes[60] or with 3-D extensions[61]. It is not yet known if these operations can extend to high rate codes, or how errors might propagate between logical gates. We also hope that an error-correction-aware graph partitioning process may improve on the partitioning scheme used in this work, possibly leading to improved error thresholds.